What is the probability of rolling a sum of a 5 or a sum of a 7?
On two 6-sided fair dice:
Here's why: Look at the dice as unique, (maybe one is red and one is blue). Then there are 36 different possible outcomes.
So there are 4 chances out of 36 (or 1 out of 9 = 11.11%) to get a sum of 5, and 6 chances out of 36 (or 1 out of 6 = 16.67%) to get a sum of 7. If you want either 5 or 7 as a sum, add the chances together to get 10 out of 36 (or 5 out of 18 = 27.78%)
What is the probability of being dealt a queen of clubs?
If only one card is dealt randomly from a deck of cards, the probability is 1/52.
The normal distribution is very difficult to work with: the probability density function is not simple [I suspect that only a few mathematics graduates from university would be able to integrate it!]. The standard normal has been tabulated in considerable detail and so, if the normal is transformed to the standard normal, most of the analyses, hypothesis testing, confidence intervals and so on can be found by looking up tables.
How do you divide standard deviation?
Standard deviation is a number and you would divide it in exactly the same way as you would divide any other number!
What is the theoretical probability of getting one head and one tail?
The theoretical probability of getting one head and one tail is 50%
What are the chances of making it into the MLB?
You definatly have to be talented to make it to the show but you always need to have a good atitude cuz u ain't gonna get in with a bad additude.
How rare is a hole in one in golf?
That depends on which hole. Shorter holes will have more common hole in ones than long ones. In general, still very rare.
What is the probability of getting 100 percent on a four question quiz?
If it is a T/F test; probability correct for each question is 0.5. Since there are 4 questions, raise 0.5 to the 4th power; e.g. (0.5)4. So, probability all correct is 0.0625. If a 4 part multiple choice, P(correct) = .25 so raise .25 to the 4th power, or .003906.
What are the chances of identical twins having twins themselves?
Women who are themselves fraternal twins have a 10% chance of producing twins, identical twin women have only a 0.6% probability of having a twin birth.
If a number cube is rolled what is the probability that the number rolled will be greater than 3?
It is 1/2 if it is a fair number cube which is rolled once.
What is the probability that the sum of the numbers on two dice is composite when they are rolled?
21/36 or 7/12 or 58.33...%
What is the difference between probability and graphs?
in probability, you only guess. while in graphs you only draw lines. but both are boring.
What are 3 numbers that have 2 or 5 factors?
3, 5 and 7 have 2 factors.
81, 625 and 2401 have 5 factors.
What is the probability that two people will have the same birthday out of 25?
Slightly more than 1 in 2.
The US penny, the symbol for one cent, is not brown... but copper. Yes, the penny is made out of copper. Copper was a very economical metal back then when it was starting to be made. However, for a short period of time, while I believe during the WW2, the penny was made out of something else, other than copper for a few years and then went back to being made out of strong copper. That is why quarters, dimes, nickels, and other coins always need to be remade, and that's why pennies are stronger and can last longer... because of the strongness of a penny. Even though it's worthless, it's pretty unworthless. Also, when copper is first exposed to air, it oxidises very quickly. Then, the common oxide of copper is brown... that's why the color of copper is brown. And therefore, the penny is brown.
Find the probability of rolling a sum of 10 with 2 dice?
Well, there are 36 options to be rolled. The only combos that will add up to ten are 5+5 and 6+4. So that's 2 options out of 36. Hence 2/36.
simplified to 1/18 OR 5.56%
Features on the probability scale?
well the question you wrote do not make any and please stop wasting your time on question like these be more specific you retard!
What is the probaility of drawing a ace from a deck of cards?
Probability is 4 aces/52 total cards = 1/13 = .07692